# A New Probe of Cosmic Birefringence Using Galaxy Polarization and Shapes

### v1

@@ Abstract
We propose a new method to search for parity-violating new physics via measurements of cosmic birefringence and demonstrate its power in detecting the topological effect originating from an axion string network with an axion-photon coupling as a motivated source of cosmic birefringence. The method, using large galaxy samples, exploits an empirical correlation between the polarization direction of the integrated radio emission from a spiral galaxy and its apparent shape. We devise unbiased minimum-variance quadratic estimators for discrete samples of galaxies with both integrated radio polarization and shape measurements. Assuming a synergy with overlapping optical imaging surveys, we forecast the sensitivity to polarization rotation of the forthcoming SKA radio continuum surveys of spiral galaxies out to $z\sim 1.5$. The angular noise power spectrum of polarization rotation using our method can be lower than that expected from CMB Stage-IV experiments, when assuming a wide survey covering $\sim 1000\,{\rm deg}^{2}$ and reaching an RMS flux of $\sim 1\,\mu{\rm Jy}$. Our method will be complementary to CMB-based methods as it will be subject to different systematics. It can be generalized to probe time-varying or redshift-varying birefringence signals.
@@
Introduction — The question of whether discrete transformations—charge conjugation (C), parity (P), and time reversal (T)—are fundamental symmetries of the Universe has led to important discoveries [1, 2, 3]. The axion was originally proposed to solve the strong CP problem in Quantum Chromodynamics [4, 5, 6, 7]. More recently, generic axions are shown to arise abundantly in string theory [8, 9]. The Chern-Simons interaction between the axion and electromagnetism leads to the parity-violating cosmic birefringence effect [10]:
@@
where $a$ is the axion field, $F^{\mu\nu}$ is the electromagnetic field strength tensor, and $\tilde{F}^{\mu\nu}\equiv(1/2)\,\epsilon_{\mu\nu\rho\sigma}\,F^{\rho\sigma}$ is its dual. As a photon travels through a background axion field, its plane of polarization rotates by an angle proportional to the net change in the axion field value along the photon’s worldline, multiplied by the axion-photon coupling constant $g_{a\gamma\gamma}$. The rotation is independent of the photon’s frequency due to the topological nature of the interaction [11, 12, 13]. Recent analyses of the cosmic microwave background (CMB) have shown hints of a rotation of its polarization anisotropies that is highly coherent across the sky, on the order of a fraction of a degree [14, 15]. Possible contamination from foregrounds is being investigated [16]. On the other hand, searches for anisotropic birefrigence have not so far uncovered a signal [17, 18, 19, 20].
@@
Measuring birefringence toward any single astrophysical source or along any single sightline is hindered by our lack of knowledge about the intrinsic polarization direction of the source. When the CMB is used as the probe, this fundamental problem is solved statistically. Rotation angles that are correlated across the sky are measurable owing to the state-of-the-art cosmological model that accurately predicts the Gaussian statistics of the primary CMB polarization anisotropies [21, 22, 23, 20]. To confidently confirm a signal of cosmic birefringence, it is important to consider alternative methods of detection based on distinct cosmological or astrophysical probes whose systematics are independent of CMB measurements.
@@
In this paper, we propose a new probe of cosmic birefringence that relies on a large sample of spiral galaxies with both radio polarization and apparent shape measurements. This is motivated by empirical findings that the integrated radio emission of nearby spiral galaxies, at least at $4.8\,\mathrm{GHz}$, is significantly polarized and aligns on average with the apparent minor axis of the optical galactic disk [24]. A plausible explanation for this alignment is that the polarized radio emission is predominantly synchrotron radiation powered by star-formation feedback and has a polarization direction perpendicular to the local toroidal magnetic field that is ordered on the galactic scale. Integrating over the entire disk results in a non-zero net polarization if the disk is inclined. It seems reasonable to assume that a similar empirical correlation is true for a cosmological population of spiral galaxies. The apparent disk shape can be measured with either optical imaging or interferometric radio imaging, with the former available from existing weak lensing catalogs [25, 26, 27] or from upcoming imaging surveys [28, 29].
@@
In reality, the polarization-shape alignment is imperfect either due to the poloidal component of the ordered magnetic field or due to randomly arising asymmetry in the disk structure. Nonetheless, one expects that the mean polarization-shape misalignment angle is zero among a sample of unrelated galaxies, as long as all the physics on the (sub-)galactic scale responsible for the polarization-shape correlation respects parity. If there is one galaxy whose integrated radio polarization is offset from the apparent minor axis to the clockwise side, it would be equally probable that another galaxy appears as its mirror reflection and has a misalignment to the counter-clockwise side (Fig. 1).
Figure 1: An illustration of the parity-violating effect cosmic birefringence has on the (mis)alignment between the linear polarization of the integrated radio emission (solid blue) and the apparent minor axis (dashed black) of a spiral galaxy. For every spiral galaxy with a mislignment angle $\Delta\theta$ (left), another galaxy whose structure is its mirror image (right) would have the opposite misalignment angle, and so $\Delta\theta$ is expected to average to zero over many galaxies. In the presence of birefringence, the integrated polarization is rotated by an angle $\alpha$ (from light to dark blue) for both the galaxy and its mirror image, and $\Delta\theta$ would be statistically biased from zero.
@@
It has been proposed that such polarization-shape correlation can be exploited to calibrate the intrinsic shapes of galaxies, assuming that polarization is not externally modified, which reduces noise in weak lensing shear measurements [30, 31, 32]. By reversing this logic, one can calibrate the intrinsic polarization direction relative to source morphology [33, 34]. With cosmic birefringence, unrelated galaxies in a patch of the sky will have their polarization rotated by nearly the same angle, leading to a polarization-shape correlation that is biased from zero.
@@
In this Letter, we construct a minimum-variance quadratic estimator for polarization rotation angles using a catalog of galaxies with both integrated radio polarization and shape measurements, and forecast the angular noise power spectrum for the estimator. Radio polarization measurements may come from the existing VLA Sky Survey [35] or the upcoming SKA continuum surveys [36], while the optical counterparts can come from an overlapping photometric galaxy catalog with shape measurements, such as from the DESI Legacy Surveys (DECaLS) [28] or the forthcoming LSST [29].
@@
Galaxy Radio Polarization — When a galaxy is viewed at an inclination, the magnetic fields appear aligned with the apparent major axis. This leads to a net linear polarization of the integrated radio emission as found in [24]. The degree of polarization of the integrated emission depends on the galaxy’s disk inclination and radio luminosity, as well as the presence of a randomly oriented component of the magnetic field throughout the galaxy. Ref. [24] found in the $4.8\,\mathrm{GHz}$ emission from nearby spiral galaxies that the net degree of polarization ranges from $\sim 1\%$ to $\sim 18\%$, with a mean of $\sim 4\%$. While the alignment between the net polarization and the minor axis is not exact in any individual galaxy, the sample of [24] shows a mean misalignment angle consistent with zero and an RMS value of $\sim 26^{\circ}$. As we will see later, the effective scatter in this misalignment angle, which limits the precision of cosmic birefringence measurements, is only as small as $\sim 5^{\circ}$, once a proper quadratic estimator is constructed. This rather remarkable result is uncovered empirically from the sample of [24], which indicates a strong correlation between the scatter in misalignment angle and disk inclination.
@@
Estimating Cosmic Birefrigence — the linear polarization of a galaxy can be described by a pair of Stokes parameters $(Q,\,U)$. A spin-2 variable $P_{\pm}$ can be defined
@@
where $I$ is the intensity  [^1] In radio astronomy, $I$ is the spectral flux density typically measured in units of jansky (Jy)., $p$ is the degree of polarization ($10<p<$ 1), and $\theta_{P}$ is the position angle of the linear polarization in the sky. The apparent elliptical shape of an inclined galaxy can be described by another spin-2 variable, the ellipticity ϵ±subscriptitalic-$\epsilon_{\pm}$ [38, p. 11]
@@
where $a$ and $b$ are the semi-minor and semi-major axes, respectively, and $\theta_{\epsilon}$ is the position angle of the semi-major axis in the sky.
Since $\theta_{P}-\theta_{\epsilon}$ is invariant under rotations of the sky, we define the following (dimensionless) spin-0 quantities
@@
where $\Delta\theta\equiv\theta_{P}-\theta_{\epsilon}-\pi/2$ is the polarization-shape misalignment angle and $|\epsilon|\equiv(a-b)/(a+b)$. The quantities $X$ and $Y$ have even and odd parity, respectively.
The quantities $p$, $|\epsilon|$, and $\Delta\theta$ are intrinsic to each galaxy. Since there is no known mechanism for macroscopic parity violation in the synchrotron radiation in the interstellar medium (ISM), we assume that $\Delta\theta$ has a distribution symmetric about $\Delta\theta=0$, and that possible parity-violating effects occur externally as a result of cosmic birefringence. Cosmic birefringence can be quantified by a sightline-dependent polarization rotation field $\alpha(\bm{\mathbf{\hat{n}}})$, coherent over some angular scale on the sky. Depending on the birefringence source, $\alpha(\bm{\mathbf{\hat{n}}})$ may be uniform on the sky [39], or behave as a Gaussian random field [40, 41], or have a highly non-Gaussian spatial pattern [13]. Thus, cosmic birefringence causes the distribution of $\Delta\theta$ to be biased from zero by an amount $\alpha(\bm{\mathbf{\hat{n}}})$ for galaxies in a local sky patch in the direction $\bm{\mathbf{\hat{n}}}$. For a first analysis, we assume that $\alpha(\bm{\mathbf{\hat{n}}})$ is independent of galaxy distance, and that $(I\,p,\Delta\theta)$ are independently drawn from the same global distribution for individual galaxies.
While the Faraday effect of the Milky Way (MW)’s ISM also rotates the polarization coherently on the sky, the rotation angle scales as the wavelength squared. This can be distinguished from the wavelength-independent birefringence effect induced by an axion [11]. We therefore assume that MW Faraday rotation is subtracted as a foreground effect prior to the following analysis. This does require observation in multiple radio bands, the effectiveness of which is subject to further study. On the other hand, Faraday effect within the source galaxy causes incoherent polarization rotation across the galactic disk, which leads to the depolarization of the integrated radio emission that is stronger at lower radio frequencies [24, 42]. This is not a concern in the following analysis, as depolarization at $4.8\,\mathrm{GHz}$ is included in the empirical polarization-shape (mis)alignment data.
Parity-odd combinations of $X$ and $Y$ have a vanishing mean over the ensemble of galaxies when $\alpha(\bm{\mathbf{\hat{n}}})=0$:
@@
In the presence of a nonzero $\alpha(\bm{\mathbf{\hat{n}}})$, the disk shape is unaffected while the integrated polarization is rotated:
@@
This implies transformation rules for $X$ and $Y$
@@
We construct an estimator for $\alpha(\bm{\mathbf{\hat{n}}})$ using $X$ and $Y$, inspired by the minimum-variance quadratic estimators for weak lensing [43] and for polarization rotation [21, 23, 22, 44] in analyzing CMB anistropies. In App. B, this estimator is confirmed by a maximum likelihood approach [45].
For CMB anisotropies, quadratic estimators for cosmic shear and for polarization rotation are orthogonal to each other to first order in each of the effects [20]. Cosmic shear induces a white-noise $Y\approx p\,|\epsilon|\,\gamma=0.002\,(p/0.1)\,(\gamma/0.02)\,|\epsilon|$, where $\sqrt{\langle\gamma^{2}\rangle}\approx 0.02$ at $z\sim 1$ [46], which however is smaller than the intrinsic $\sigma_{Y}=0.004$ (Eq. (13)). With both weak lensing and galaxy intrinsic alignment, a correlated noise contribution to $\hat{\alpha}$ arises, but with a magnitude significantly smaller than the noise floor derived here. We will therefore disregard weak lensing in the following calculation.
To estimate the polarization rotation $\alpha(\bm{\mathbf{\hat{n}}})$ using a single galaxy in the direction $\bm{\mathbf{\hat{n}}}$, we write down an estimator $\hat{\alpha}(\bm{\mathbf{\hat{n}}})$ as the general combination of the observed quantities $\tilde{X}$ and $\tilde{Y}$ (including measurement noise) for this galaxy up to quadratic terms:
where the coefficients $u,v,u^{\prime},v^{\prime},w$ are to be determined. We omit the dependence on $\bm{\mathbf{\hat{n}}}$ for clarity of notation whenever only a single galaxy is concerned.
Following the idea underlying the minimum-variance quadratic estimator, we require that $\hat{\alpha}$ be an unbiased estimator of $\alpha$. That is, $\expectationvalue{\hat{\alpha}}=\alpha$ when averaged over an ensemble of galaxies with independent intrinsic properties. Subject to this requirement, we then minimize the variance of $\hat{\alpha}$ under the null hypothesis $\alpha=0$. To carry out the calculation analytically, we approximate $X$ as a Gaussian random variable with mean $\overline{X}$ and variance $\sigma_{X}^{2}$, and $Y$ a Gaussian random variable with zero mean and variance $\sigma_{Y}^{2}$. Since $X$ and $Y$ are only approximately Gaussian, this gives a nearly optimal quadratic estimator. The true optimal estimator could be obtained by numerically minimizing $\expectationvalue{\hat{\alpha}^{2}}$, given the empirically measured distributions for $Ip$, $|\epsilon|$, and $\Delta\theta$.
Observed quantities have measurement uncertainties. If the radio intensity measurement has an RMS uncertainty of $\sigma_{I}$, then the uncertainty on the polarization degree $p$ is $\sigma_{I}/I$, which equals the inverse unpolarized signal-to-noise ratio (SNR) of the galaxy. We find that with realistic deep optical imaging surveys that include shape measurement, the radio polarization measurement error dominates over the error in ellipticity measurement. This leads to the overall errors for $\tilde{X}$ and $\tilde{Y}$, also assumed to be Gaussian:
$\tilde{\sigma}_{X,Y}^{2}=\sigma_{X,Y}^{2}+(1/2)\,\left(\Delta\epsilon/\mathrm{SNR}\right)^{2},$ (10)
where $\Delta\epsilon\approx 0.3$ is the intrinsic spread in galaxy ellipticity, and $\sigma^{2}_{X,Y}$ are the intrinsic variances of $X$ and $Y$.
Since large rotation angles have been ruled out by observation, it is reasonable to assume that polarization rotation is small $\alpha\ll 1$. We find the following unbiased minimum-variance quadratic estimator (detailed derivation can be found in App. A):
@@
Conceptually, we have obtained the expected result that the estimator for the parity-violating $\alpha(\bm{\mathbf{\hat{n}}})$ is constructed from the parity-odd observables $Y$ and $X\,Y$ only. Since $\expectationvalue{\hat{\alpha}(\bm{\mathbf{\hat{n}}})}=0$ when $\alpha(\bm{\mathbf{\hat{n}}})=0$, a statistically significant deviation of $\expectationvalue{\hat{\alpha}(\bm{\mathbf{\hat{n}}})}$ from zero toward any direction $\bm{\mathbf{\hat{n}}}$ over a number of galaxies would be evidence for cosmic birefringence in that direction.
Sensitivity — From the observables $p$, $|\epsilon|$, and $\Delta\theta$, $X$ and $Y$ for each galaxy can be computed. We obtain values for $p$ and $\Delta\theta$ for 13 nearby spiral galaxies from Ref. [24], and compute values for $|\epsilon|$ from the angular major and minor axes given in the SIMBAD Astronomical Database [47]. We exclude M31 from that sample, as its close proximity and its particularly high degree of polarization at $18\%$ hint at possible systematics. This results in the following:
@@
Despite the small sample size used here, we find $\sqrt{\langle|\epsilon|^{2}\rangle}=0.30$, in good agreement with the typical value for cosmological samples. It is more uncertain whether the properties of the integrated radio polarization of this sample are representative of a cosmological sample. Verification would require analyses of samples of distant galaxies (e.g. [35, 48, 42]). Using values in Eq. (13) for $\overline{X}$, $\sigma_{X}$ and $\sigma_{Y}$ while fixing $\overline{Y}=0$, Eq. (12) gives the standard error in the rotation estimator for a single galaxy, $\sqrt{\expectationvalue{\hat{\alpha}^{2}}}=5^{\circ}$, for negligible measurement error.
@@
Remarkably, $\sqrt{\expectationvalue{\hat{\alpha}^{2}}}=5^{\circ}$ is about five times smaller than the standard deviation of $\Delta\theta$ at $26^{\circ}$ as estimated from nearby spiral galaxies. We attribute this startling difference to an anti-correlation between the scatter in the misalignment angle $\Delta\theta$ and the product $p\,|\epsilon|$. According to Eq. (4), when computing the statistics of $X$ and $Y$ in Eq. (13), galaxies with large $\Delta\theta$ values are weighted less due to smaller $p\,|\epsilon|$ values.
@@
This anti-correlation has a geometrical origin (the effect is quantitatively modelled in [24]). The intrinsic misalignment angle has a large variance for galaxies showing a low integrated polarization degree and a low ellipticity, as these galaxies mostly have a low inclination and appear rather circular. Since $p$ and $|\epsilon|$ both increase with inclination, the prefactor $p\,|\epsilon|$ effectively down-weights face-on galaxies and emphasizes edge-on galaxies on which birefringence would have a larger detectable effect. While we do not pursue it in this work, it seems possible to us that if this anti-correlation is accurately modelled as a function of the inclination, then an improved definition of $X$ and $Y$ could be derived to replace Eq. (4).
To make sure that the surprisingly small value of $\sqrt{\langle\hat{\alpha}^{2}\rangle}=5^{\circ}$ is not a fluke of low-number statistics from only 13 galaxies, we perform bootstrapping on the data extracted from Refs. [24, 47]. In each run, 4,000 pairs of $(p|\epsilon|,\Delta\theta)$ are generated by independent draws from the 13 galaxies with replacement. The value of $\expectationvalue{\hat{\alpha}^{2}}$ is then computed from the $4,000$ artificial pairs using Eq. (12). We repeat this procedure 2,000 times, and obtain $\sqrt{\expectationvalue{\hat{\alpha}^{2}}}=(44\pm 5)^{\circ}$ (see Fig. 2), strongly suggesting that the anti-correlation between $\Delta\theta$ and $p\,|\epsilon|$ is real.
@@
Figure 2: Histogram of the standard error of $\hat{\alpha}$ when the anti-correlation between $p|\epsilon|$ and $\Delta\theta$ is erased by bootstrapping using a sample of 13 galaxies, repeated 2,000 times. The dashed line marks the standard error of $\hat{\alpha}$ measured from the 13 galaxies without boostrapping. Its significantly lower value compared to the bootstrapped distribution indicates that there is a genuine anti-correlation between $\Delta\theta$ and $p\,|\epsilon|$.
@@
Galaxy Survey — Under the flat-sky approximation, the estimator is given by a finite sum over $N$ discrete galaxies
@@
where $\Omega$ is survey area in steradians. The variance of the estimator is then
@@
In the absence of birefringence, the estimation for the rotation power spectrum is therefore limited by a white-noise floor
$C_{\hat{\alpha}\hat{\alpha}}(\mathbf{L})=(\Omega/N)\,\expectationvalue{\hat{\alpha}^{2}},$ (17)
analogous to the shot noise in cosmic shear estimation from a discrete sample of galaxies.
We adapt the procedure used in Ref. [49] to count the number of available spiral galaxies. We first compute the redshift-dependent radio luminosity function of spiral galaxies. The radio luminosity function at $4.8\,\mathrm{GHz}$ at redshift $z$ can be expressed in terms of the radio luminosity function at $1.4\,\mathrm{GHz}$ at $z=0$, taking into account luminosity evolution [50] and the spectral index ($S_{\nu}\propto\nu^{\alpha}$ where $\alpha=-0.7$ [51, 52]). We choose $4.8\,\mathrm{GHz}$ because Ref. [24] established the empirical correlation at this frequency and because Faraday rotation in both the source and the MW is greatly reduced at higher frequencies. Next, the local radio luminosity function can be obtained from the far-infrared (FIR) luminosity function [53, 54] by exploiting an approximately linear relation between the radio and FIR luminosities of star-forming galaxies [55, 56].
The number of spiral galaxies with an observed spectral flux density $S_{\nu}$ above a threshold $S_{\nu}^{\mathrm{min}}$ is given by

### v2

@@ Abstract
We propose a novel statistical method to measure cosmic birefringence and demonstrate its power in probing parity violation due to axions. Exploiting an empirical correlation between the integrated radio polarization direction of a spiral galaxy and its apparent shape, we devise an unbiased minimum-variance estimator for the rotation angle, which should achieve an uncertainty of $5^{\circ}$$15^{\circ}$ per galaxy. Large galaxy samples from the forthcoming SKA continuum surveys, together with optical shape catalogs, promise a comparable or even lower noise power spectrum for the rotation angle than in the CMB Stage-IV (CMB-S4) experiment, with different systematics.
@@
Introduction — The study of fundamental discrete symmetries have led to important discoveries [1, 2, 3]. The axion was originally proposed to solve the strong CP problem in Quantum Chromodynamics [4, 5, 6, 7]. Generic axions were later shown to arise abundantly in string theory [8, 9]. The Chern-Simons interaction between the axion and electromagnetism causes the parity-violating cosmic birefringence effect [10]:
@@
where $a$ is the axion field, $F^{\mu\nu}$ is the field strength tensor, and $\tilde{F}^{\mu\nu}$ its dual. As a photon travels, its polarization plane rotates by an angle proportional to the net change in the unwrapped axion field value along its worldline, multiplied by the axion-photon coupling $g_{a\gamma\gamma}$ [11, 12, 13]. Recent analyses of the cosmic microwave background (CMB) have shown hints of birefringence uniform across the sky at a fraction of a degree [14, 15]. Possible foreground contamination is being investigated [16]. On the other hand, searches for anisotropic birefringence have not uncovered a signal [17, 18, 19, 20].
@@
Measuring birefringence toward any single astrophysical source is hindered by our lack of knowledge about the intrinsic polarization direction. The use of the CMB as a probe statistically solves this problem. Rotation angles correlated across the sky are measurable following accurate cosmological prediction for the Gaussian statistics of the primary CMB anisotropies [21, 22, 23, 20]. To robustly confirm cosmic birefringence, it is important to consider alternative detection methods based on independent cosmological or astrophysical probes.
@@
In this paper, we propose a new probe of cosmic birefringence that relies on surveying spiral galaxies with both radio polarization and apparent shape measurements. This is motivated by empirical findings that the integrated radio emission of nearby spiral galaxies at $4.8\,\mathrm{GHz}$ is significantly polarized and aligns on average with the apparent minor axis of the optical galactic disk [24]. A plausible explanation for this alignment is that the polarized radio emission is predominantly synchrotron radiation powered by star formation feedback and has a polarization direction locally perpendicular to the ordered toroidal magnetic field. Integrating over an inclined disk results in a non-zero net polarization. It seems reasonable to assume that such correlation is true for a cosmological population of spirals. The shape ellipse can be measured with either optical or interferometric radio imaging, with the former available from weak lensing catalogs [25, 26, 27] or from imaging surveys [28, 29].
@@
The polarization-shape alignment is imperfect for each single galaxy. Nonetheless, one expects that the mean misalignment angle is zero among unrelated galaxies, as long as all relevant physics on the (sub-)galactic scale respects parity. If there is one galaxy whose integrated polarization is offset from the apparent minor axis to the clockwise side, it should be equally probable that another galaxy appears as its mirror reflection and has a counter-clockwise misalignment (Fig. 1).
Figure 1: (Mis)alignment between the polarization of the integrated radio emission (solid blue) and the apparent minor axis (dashed black) of a spiral galaxy.
@@
It has been proposed that polarization-shape correlation can be exploited to calibrate galaxy intrinsic shapes and reduce noise in weak lensing shear measurements [30, 31, 32, 33]. Reversing this logic, one can calibrate the intrinsic polarization direction to source morphology [34, 35]. With cosmic birefringence, unrelated galaxies in a patch of the sky will have polarization rotated by nearly the same angle, leading to a mean misalignment biased from zero.
@@
In this Letter, we construct a minimum-variance quadratic estimator for polarization rotation using galaxies with both integrated radio polarization and shape measurements, and forecast its angular noise power spectrum. Radio polarization measurements may come from the VLA Sky Survey [36] or the upcoming SKA continuum surveys [37], while the optical counterparts can come from photometric galaxy catalogs with shape measurements, such as from the DESI Legacy Surveys (DECaLS) [28] or LSST [29].
@@
Galaxy Radio Polarization — The polarization degree of the integrated radio emission depends on the disk inclination and radio luminosity, as well as the randomly oriented component of the magnetic field throughout the galaxy. Ref. [24] found that the net polarization degree of nearby spirals at $4.8\,\mathrm{GHz}$ is $\sim 1\%$$18\%$ with a mean $\sim 4\%$. The sample of [24] shows a mean misalignment angle consistent with zero and an RMS value of $\sim 26^{\circ}$. As we will show, the effective scatter in this misalignment angle, which limits the precision of birefringence measurements, can be reduced to $5^{\circ}$$15^{\circ}$, once a proper estimator is used.
@@
Estimating Cosmic Birefrigence — Linear polarization is described by the Stokes parameters $(Q,\,U)$. Define a spin-2 variable:
@@
where $I$ is the intensity  [^1] In radio astronomy, $I$ is the spectral flux density typically measured in units of jansky (Jy)., $p$ is the polarization degree ($10<p<$ 1), and $\theta_{P}$ is the polarization position angle. The apparent elliptical shape of an inclined galaxy can be described by the spin-2 ellipticity [39, p. 11]
@@
where $a$ and $b$ are the semi-minor and semi-major axes, respectively, and $\theta_{\epsilon}$ is the position angle of the semi-major axis. Since $\theta_{P}-\theta_{\epsilon}$ is rotationally invariant, we define dimensionless spin-0 quantities
@@
where $\Delta\theta\equiv\theta_{P}-\theta_{\epsilon}-\pi/2$ is the misalignment angle and $|\epsilon|\equiv(a-b)/(a+b)$. The quantities $X$ and $Y$ have even and odd parity, respectively.
The quantities $p$, $|\epsilon|$, and $\Delta\theta$ are intrinsic to each galaxy. Since there is no known evidence for macroscopic parity violation in the interstellar medium (ISM) correlated between many galaxies over intergalactic distances, we assume that $\Delta\theta$ has a distribution symmetric about zero, and that correlated parity-violating effects only occur externally as a result of cosmic birefringence, which can be quantified by a polarization rotation field $\alpha(\bm{\mathbf{\hat{n}}})$ coherent over some scales on the sky. This field may be uniform on the sky [40], or behave as a Gaussian random field [41, 42], or is highly non-Gaussian [13]. Thus, cosmic birefringence causes the distribution of $\Delta\theta$ to be biased from zero by an amount $\alpha(\bm{\mathbf{\hat{n}}})$ for galaxies in a local sky patch in the direction $\bm{\mathbf{\hat{n}}}$. For a first analysis, we assume that $\alpha(\bm{\mathbf{\hat{n}}})$ is independent of galaxy distance, and that $(I\,p,\Delta\theta)$ are independently drawn from the same global distribution for individual galaxies.
The Faraday effect in the Milky Way (MW)’s ISM also rotates the polarization coherently on the sky, with an angle that scales with the wavelength squared. This can be distinguished from the wavelength-independent effect of the axion [11]. We assume that MW Faraday rotation is corrected prior to the following analysis. This does require observation in multiple radio bands, the effectiveness of which is subject to further study. On the other hand, inhomogeneous Faraday rotation within the source galaxy causes depolarization of the integrated emission particularly at lower radio frequencies [24, 43]. This is not a concern here, as depolarization at $4.8\,\mathrm{GHz}$ is included in the polarization-shape (mis)alignment data.
Parity-odd combinations of $X$ and $Y$ have a vanishing mean over the galaxy ensemble when $\alpha(\bm{\mathbf{\hat{n}}})=0$:
@@
With a non-zero $\alpha(\bm{\mathbf{\hat{n}}})$, the polarization is rotated while the disk shape is unaffected:
@@
This implies for $X$ and $Y$
@@
We construct an unbiased, minimum-variance quadratic estimator for $\alpha(\bm{\mathbf{\hat{n}}})$ using $X$ and $Y$ following weak-lensing  [44] and birefringence [21, 23, 22, 45] applications. It can be derived via maximizing the likelihood [46] (App. B).
For a single galaxy, it is
@@
The errors in $\tilde{X}$ and $\tilde{Y}$ including measurement noises are:
$\tilde{\sigma}_{X,Y}^{2}=\sigma_{X,Y}^{2}+f\,\left(\sigma_{\epsilon}/\mathrm{SNR}\right)^{2},$ (11)
where $\sigma_{\epsilon}\equiv\sqrt{\expectationvalue{|\epsilon|^{2}}}\approx 0.3$ is the intrinsic galaxy shape error, and $\sigma^{2}_{X,Y}$ are the intrinsic variances of $X$ and $Y$. Radio interferometry with pure noise has $f=1$ (detailed derivation in App. A). In the following, we derive numerical forecast assuming the conservative $f=1$.
Sensitivity — From the observed $p$, $|\epsilon|$, and $\Delta\theta$, $X$ and $Y$ for each galaxy can be computed. We obtain values for $p$ and $\Delta\theta$ for 13 nearby spirals from Ref. [24], and compute values for $|\epsilon|$ using SIMBAD shapes [47]. We exclude M31 as its proximity and high degree of polarization ($18\%$) hint at systematics. This results in the following:
@@
Despite the small sample size, we find $\sqrt{\langle|\epsilon|^{2}\rangle}=0.30$, consistent with what is assumed in galaxy weak lensing. It is more uncertain whether the radio properties of this sample are representative of a cosmological sample. Verification would require analyzing distant galaxies (e.g. [36, 48, 43]). Using values in Eq. (12) for $\overline{X}$, $\sigma_{X}$ and $\sigma_{Y}$ and setting $\overline{Y}=0$, Eq. (9) gives $\langle\hat{\alpha}\rangle=0.5^{\circ}$ and a sample standard deviation $4.4^{\circ}$, consistent with no birefringence; Eq. (10) forecasts $\sqrt{\expectationvalue{\hat{\alpha}^{2}}}=5^{\circ}$ per galaxy in the limit of negligible measurement error.
@@
Remarkably, $\sqrt{\expectationvalue{\hat{\alpha}^{2}}}=5^{\circ}$ is five times smaller than the standard deviation of $\Delta\theta$ at $26^{\circ}$ as estimated from nearby spirals. We attribute this improvement to an anti-correlation between the scatter in the misalignment angle $\Delta\theta$ and the product $p\,|\epsilon|$. According to Eq. (4), when computing the statistics of $X$ and $Y$ in Eq. (12), galaxies with large $\Delta\theta$ are down-weighted due to smaller $p\,|\epsilon|$ values.
@@
This anti-correlation has a geometrical origin [24]. The intrinsic misalignment angle has a large variance for galaxies showing a low $p$ and a low $|\epsilon|$, as these low-inclination galaxies appear rather circular. Since $p$ and $|\epsilon|$ both increase with inclination, $p\,|\epsilon|$ down-weights face-on galaxies and emphasizes edge-on ones on which birefringence would be more detectable. While we do not pursue it in this work, it seems possible to accurately model this anti-correlation as a function of the inclination and derive an improved definition of $X$ and $Y$ to replace Eq. (4).
The $5^{\circ}$ error may result from a particularly good polarization-shape alignment in the small galaxy sample. Larger samples are required to refine this figure. Our primary claim is not this precise error value, but rather that our novel estimator effectively exploits the inclination-dependent statistics of the misalignment angle and has the potential to dramatically improve sensitivity to birefringence.
Fig. 2 estimates the uncertainty in this number by bootstrapping and demonstrates that the anti-correlation between $p|\epsilon|$ and $\Delta\theta$ is not a statistical fluke. In each run, $p|\epsilon|$ and $\Delta\theta$ are independently drawn with replacement from the Stil et al. sample to form 13 artificial galaxies, from which $\expectationvalue{\hat{\alpha}^{2}}$ is computed. In 2,000 runs, 97.5% produce a higher estimator uncertainty than $5^{\circ}$, strongly supporting a genuine anti-correlation between $\Delta\theta$ and $p\,|\epsilon|$. We note that $\sqrt{\langle\hat{\alpha}^{2}\rangle}=13^{\circ}$ if NGC 4565, an edge-on galaxy with $p=10.5\%$, is excluded. Its ellipticity implies a $\sim 10\%$ chance of occurrence for such “good” galaxies. Therefore, it may be expected that there is one such galaxy out of a sample of 13.
@@
Figure 2: Distribution of the standard deviation of $\hat{\alpha}$ from 2000 bootstrap sets of 13 galaxies (blue), compared to the case of erasing the anti-correlation between $p|\epsilon|$ and $\Delta\theta$ (orange). The dashed line marks the standard deviation of $\hat{\alpha}$ measured from the Stil et al. sample of 13 galaxies. Its lower value compared to the orange distribution indicates a genuine anti-correlation between $\Delta\theta$ and $p\,|\epsilon|$.
@@
Galaxy Survey — Under the flat-sky approximation, the estimator is given by a sum over $N$ galaxies
@@
where $\Omega$ is survey area in steradians. The estimator has a variance
@@
Without birefringence, the rotation estimator has a white-noise angular power spectrum
$C_{\hat{\alpha}\hat{\alpha}}(\mathbf{L})=(\Omega/N)\,\expectationvalue{\hat{\alpha}^{2}}.$ (16)
We adapt the procedure used in Ref. [49] to count available spiral galaxies. We first compute the redshift-dependent radio luminosity function of spiral galaxies. The radio luminosity function at $4.8\,\mathrm{GHz}$ at redshift $z$ can be related to that at $1.4\,\mathrm{GHz}$ at $z=0$, taking into account luminosity evolution [50] and the spectral index ($S_{\nu}\propto\nu^{\alpha}$ where $\alpha=-0.7$ [51, 52]). We choose $4.8\,\mathrm{GHz}$ because Ref. [24] established the empirical correlation at this frequency and Faraday rotation in both the source and the MW is greatly reduced at higher frequencies. Next, the local radio luminosity function can be obtained from the far-infrared (FIR) luminosity function [53, 54] using a nearly linear relation between the radio and FIR luminosities of star-forming galaxies [55, 56, 57, 58, 59, 60].
The number of spirals with an observed spectral flux density $S_{\nu}$ above a threshold $S_{\nu}^{\mathrm{min}}$ is